find the x and y intercepts of 3x+5y=30
Answer by sdmmadam@yahoo.com(530)

find the x and y intercepts of 3x+5y=30
3x+5y=30 ----(1)
Dividing by 30
3x/30 + 5y/30 = 30/30
x/10+y/6 = 1 ----(2)
(2) is in the standard intercept form
x/a + y/b = 1 ----(*)
where a = x - intercept and b = y-intercept
Therefore comparing (2) with (*)
we have x -intercept = 10
and y - intercept = 6
Note:The line (1) cuts the x -axis at (10,0) and
the y-axis at (0,6)